# Evolutionary Dynamics of Acid Resistance in Tumors: A Mathematical Model

### v1

@@ Abstract
Acidosis in tumors arises from reprogrammed metabolism and compromised vasculature, creating a harsh, acidic microenvironment that drives the evolutionary selection of acid-resistant cell phenotypes. A mathematical model is proposed to integrate phenotypic evolution, microenvironmental acidification, and tumor density dynamics. Three key mechanisms are incorporated in it: frequency-dependent selection favoring acid-resistant cells below a critical pH, stress-induced phenotypic switching, and a positive feedback loop where resistant cells produce excess acid that intensifies selection pressure. The well-posedness of the model is established. Through numerical simulations across biologically relevant parameter regimes, we identify two therapeutically targetable parameters: baseline acid clearance rate (a proxy for vascular perfusion) and a protection factor (representing acid-resistance ”machinery” effectiveness) as critical bifurcation parameters for resistance evolution. The model exhibits qualitatively distinct dynamics depending on phenotypic plasticity levels. In low-plasticity tumors, both parameters exhibit sharp bifurcations with strong parameter interactions: clearance and protection effects are context-dependent, with therapeutic interventions effective only within specific parameter ranges. In high-plasticity tumors, both parameters produce continuous, monotonic responses with independent, additive effects. These regime-dependent dynamics suggest that treatment strategies should adapt to tumor plasticity: in the formar, targeting perfusion alone is typically sufficient, though sequential therapy may be required if the perfusion rate approaches or exceeds the bifurcation threshold, whereas in the latter treatment might benefit from combination therapies addressing both parameters simultaneously.
@@ 1 Introduction
Tumors thrive in a complex microenvironment, distinguishing them from regular tissue through multiple physiological changes. One of the most common features is extracellular acidosis, as tumor pH values usually range from 6.5 to 7.0, while normal tissue maintains a pH of 7.4 [1]. This phenomenon is the result of two related processes: increased glycolytic metabolism in cancer cells, which is known as the Warburg effect [2], and irregular tumor vasculature that inhibits the clearance of waste products, including protons (H+ ions) [3].
@@
Despite increasing recognition of the impact of tumor acidosis and the evolution of acid-resistant cells, mathematical models that incorporate these dynamics into a tumor growth model remain scarce. Mathematical modeling offers a convenient approach to capture these complex processes and to help predict how therapeutic interventions could mitigate them [13, 14]. Previous models have examined the impact of the Warburg effect on the metabolic state and acidification of a tumor [15], the evolutionary dynamics of tumor heterogeneity [16], and the spatial distribution of pH in tumors [17, 18], but none seem to have incorporated phenotypic plasticity, frequency-dependent selection, and positive feedback through increasing environmental acidification within a unified framework.
In this work, we develop a mathematical model that integrates these mechanisms to investigate how two key parameters, baseline acid clearance and acid-resistance “machinery” effectiveness, affect resistance evolution. Through bifurcation analysis across different levels of phenotypic plasticity, we identify regime-dependent therapeutic strategies and establish plasticity as a critical determinant of treatment response in acidic tumor microenvironments.
@@ 2.1 State Variables
Clinically relevant tumor pH values typically fall below physiological levels, reflecting varying degrees of acidosis. For instance, a mild decrease in pH to 7.2 corresponds to $h\approx 1.6$, moderate acidosis at pH 7.0 yields $h\approx 2.5$, and more severe acidification at pH 6.8 and 6.5 correspond to $h\approx 4.0$ and $h\approx 7.9$, respectively. These values delineate the physiological range over which the tumor microenvironment exerts selective pressure favoring acid-resistant phenotypes.
@@ 2.5 Equilibrium Points
Concerning the stability of $E_{0}$, we have the following result based on the basic reproduction number of the total population.
Let
$\mathcal{R}_{0}=\displaystyle\frac{r}{D(h,x)}$ (18)
denote the basic reproduction number of the total cell population of the tumor at acidity level $h$ and fraction of resistant cells $x$.
@@
We were unable to obtain stability results for this equilibrium, and we were also unable to compute the positive equilibria with $\nu>0$, corresponding to $0<x^{*}<1$.
2.6 Parameter Value Estimation
@@ 2.6.4 Acid Production and Clearance
$\gamma_{0}$ $\in[0.02,0.05]\text{ day}^{-1}$ (maximal unit clearance rate) (39)
@@ 2.7 Baseline Parameter Set
$\gamma_{0}$ Maximal vascular clearance unit rate day-1 0.03 [0.02, 0.05]
@@ 3.1 Methods
We explored two distinct parameter regimes, both biologically plausible but differing in phenotypic plasticity and resistance effectiveness.
@@ 3.2.3 Bifurcation with Respect to Protection Factor
Intermediate clearance ($\gamma_{0}=0.039$ day-1): At the critical clearance rate, protection factor produces a sharp bifurcation. Above $\varphi\approx 0.58$, the system equilibrates to high resistance ($x^{*}>0.95$), while below this threshold, resistance is almost suppressed ($x^{*}<0.1$). This indicates that at marginal perfusion levels, acid-resistance machinery effectiveness becomes a critical determinant of evolutionary outcomes.
@@
Figure 3: Regime A: Context-dependent bifurcation with respect to protection factor. Equilibrium resistant fraction $x^{*}$ as a function of protection factor, $\varphi$, at three clearance rates: (Left) $\gamma_{0}=0.025$ day-1 (below bifurcation threshold), (Middle) $\gamma_{0}=0.039$ day-1 (at bifurcation threshold), (Right) $\gamma_{0}=0.052$ day-1 (above bifurcation threshold). Protection factor produces a sharp bifurcation only at intermediate clearance rates, demonstrating parameter interaction. At low clearance, resistance is inevitable regardless of $\varphi$; at high clearance, sensitivity is maintained regardless of $\varphi$. This indicates that targeting acid-resistance machinery is only effective within a specific range of vascular perfusion levels.
@@ 4 Discussion and Conclusions
Our model’s identification of plasticity as a critical determinant of treatment response emphasizes the need for biomarkers distinguishing these tumor phenotypes. Potential markers of phenotypic plasticity include expression levels of epigenetic modifiers (EZH2, DNMT1), epithelial-mesenchymal transition markers, or functional assays measuring adaptation rates to environmental stress [9, 10, 31]. This framework provides a foundation for regime dependent interventions where treatment strategies are tailored to the tumor’s evolutionary capacity: perfusion-focused therapy for low-plasticity tumors versus combination approaches for high-plasticity tumors.

### v2

@@ Abstract
Acidosis in tumors arises from reprogrammed metabolism and compromised vasculature, creating a harsh, acidic microenvironment that drives the evolutionary selection of acid-resistant cell phenotypes. A mathematical model is proposed to integrate phenotypic evolution, microenvironmental acidification, and tumor density dynamics. Three key mechanisms are incorporated in it: frequency-dependent selection favoring acid-resistant cells below a critical pH, stress-induced phenotypic switching, and a positive feedback loop where resistant cells produce excess acid that intensifies selection pressure. The well-posedness of the model is established. Through numerical simulations across biologically relevant parameter regimes, we identify two therapeutically targetable parameters: baseline acid clearance rate (a proxy for vascular perfusion) and a protection factor (representing acid-resistance ”machinery” effectiveness) as critical bifurcation parameters for resistance evolution. The model exhibits qualitatively distinct dynamics depending on phenotypic plasticity levels. In low-plasticity tumors, both parameters exhibit sharp bifurcations with strong parameter interactions: clearance and protection effects are context-dependent, with therapeutic interventions effective only within specific parameter ranges. In high-plasticity tumors, both parameters produce continuous, monotonic responses with independent, additive effects. These regime-dependent dynamics suggest that treatment strategies should adapt to tumor plasticity: in the formar, targeting perfusion alone is typically sufficient, though sequential therapy may be required if the perfusion rate approaches or exceeds the bifurcation threshold, whereas in the latter treatment might benefit from combination therapies addressing both parameters simultaneously.{These findings suggest that a low-dimensional model can identify therapeutically targetable parameters governing resistance evolution, suggesting interventions that may prevent or reverse the harmful effect of acid-resistant phenotypes, which are associated with chemotherapy failure, immune evasion, and metastatic progression.
@@ 1 Introduction
Tumors thrive in a complex microenvironment, distinguishing them from regular tissue through multiple physiological changes. One of the most common features is extracellular acidosis, as tumor pH values usually range from 6.5 to 7.0, while normal tissue maintains a pH of 7.4 [1]. This phenomenon is the result of two related processes: increased glycolytic metabolism in cancer cells, which is known as the Warburg effect [2], and irregular tumor vasculature that inhibits the clearance of waste products, including protons (H+ ions) [3]. This acid-resistant phenotype has been documented across multiple cancer types, including colorectal cancer, where acidosis promotes DNA damage responses and genomic instability [4], and lung cancer, where it facilitates metastatic colonization through extracellular matrix remodeling [5]. Clinically, acidosis-driven resistance has been implicated in treatment failure across multiple modalities: acidic extracellular pH reduces uptake of weakly basic chemotherapeutics such as doxorubicin [6], impairs T-cell cytotoxicity and checkpoint inhibitor efficacy [7], and promotes survival of radioresistant cell populations [8].
@@
Several mathematical frameworks have addressed some components of this problem. Game-theoretic models have captured frequency-dependent selection among tumor phenotypes: Archetti [18] framed glycolysis as a collective action problem where acid production serves as a public good, while Kaznatcheev et al. [19] extended this to a “double goods” game incorporating both acidification and vascularization, demonstrating how polyclonal equilibria emerge from replicator dynamics. More recently, Fiandaca et al. [20] developed a partial integro-differential equation model coupling continuous phenotypic evolution with oxygen and lactate dynamics, showing how spatial gradients drive selection for acid-resistant traits. However, these models treat phenotypes as either fixed strategies or continuously evolving traits, omitting stress-induced phenotypic switching, the rapid, reversible transitions triggered by environmental stress. Furthermore, none incorporate a positive feedback loop where acid-resistant cells, through upregulated glycolysis associated with proton export machinery, produce excess acid that intensifies the very selection pressure favoring resistance. Our model addresses this gap by integrating frequency-dependent selection, stress-induced plasticity, and resistance-driven acidification within a unified ODE framework.
In this work, we develop a mathematical model that integrates these mechanisms to investigate how two key parameters, baseline acid clearance and acid-resistance “machinery” effectiveness, affect resistance evolution. Through bifurcation analysis across different levels of phenotypic plasticity, we identify regime-dependent therapeutic strategies and establish plasticity as a critical determinant of treatment response in acidic tumor microenvironments. Specifically, we ask: can a low-dimensional ODE model identify therapeutically targetable parameters that govern resistance evolution, and thereby suggest interventions capable of preventing or reversing the tumor phenotypes associated with treatment failure, immune evasion, and metastatic progression?
@@ 2.1 State Variables
Clinically relevant tumor pH values typically fall below physiological levels, reflecting varying degrees of acidosis. For instance, a mild decrease in pH to 7.2 corresponds to $h\approx 1.6$, moderate acidosis at pH 7.0 yields $h\approx 2.5$, and more severe acidification at pH 6.8 and 6.5 correspond to $h\approx 4.0$ and $h\approx 7.9$, respectively. These values delineate the physiological range over which the tumor microenvironment exerts selective pressure favoring acid-resistant phenotypes. While $h$ is mathematically defined for all positive values (as established in Theorem 1), biologically relevant tumor microenvironments constrain $h$ to approximately $[1,8]$, corresponding to the pH range $[6.5,7.4]$ [1]. Values beyond this range represent extreme conditions: $h<1$ corresponds to alkalosis (pH >7.4>7.4), rarely observed in solid tumors, while $h>8$ (pH <6.5<6.5) induces extensive cell death that naturally limits further acidification [21, 22]. Rather than explicitly modeling extracellular buffering capacity, our model incorporates implicit saturation mechanisms that prevent unbounded acid accumulation: the glycolytic inhibition factor $G(h)$ reduces acid production as $h$ increases, while the Hill function $d_{S}(h)$ causes population decline through acid-induced mortality. These mechanisms, combined with density-dependent clearance $\Gamma(n)$, ensure $h$ remains bounded within the physiologically relevant range in all simulations.
@@ 2.5 Equilibrium Points
Note that, for $h\equiv 1$ and $x\equiv x^{*}$, we can solve the equation for the total cell population density, $n$, in (4) explicitly, using the notation $D^{*}=D(1,x^{*})$:
$n(t)=\displaystyle\frac{D^{*}-r}{\left(r+\frac{D^{*}-r}{n(0)}\right)e^{(D^{*}-r)T}-r},$ (18)
whereby
$\lim_{t\to\infty}n(t)=\left\{\begin{aligned} 0,&\quad\mbox{if\quad}\frac{r}{D^{*}}<1,\\ 1-\frac{1}{D^{*}/r},&\quad\mbox{if\quad}\frac{r}{D^{*}}>1.\end{aligned}\right$. (19)
We can see in (19) that
$\mathcal{R}_{0}=\displaystyle\frac{r}{D^{*}}$ (20)
is the basic reproduction number of the total cell population of the tumor at acidity level $h=1$ and fraction of resistant cells $x=x^{*}$. Concerning the stability of $E_{0}$, we have the following result based on this basic reproduction number.
@@
This theorem suggests that $\mathcal{R}_{0}$ acts as a bifurcation parameter and, at $\mathcal{R}_{0}=1$ we have a forward bifurcation from the local stability of $E_{0}$ for $\mathcal{R}_{0}<1$ to its loss of stability for $\mathcal{R}_{0}>1$, when another (positive) equilibrium appears and inherits the local stability.
@@
We were unable to prove stability results for this equilibrium, and we were also unable to find an explicit formula for the positive equilibria with $\nu>0$, corresponding to $0<x^{*}<1$.
2.6 Parameter Estimation
@@ 2.6.4 Acid Production and Clearance
$\gamma_{0}$ [ $\in[\ {0.01,0.06}]\text{ day}^{-1}$ (maximal unit clearance rate) (41)
@@ 2.7 Baseline Parameter Set
$\gamma_{0}$ Maximal vascular clearance unit rate day-1 0.03 [ 0.01, 0.06]
@@ 3.1 Methods
Parameter Regimes: We explored two distinct parameter regimes, both biologically plausible but differing in phenotypic plasticity and resistance effectiveness. The specific parameter values defining these regimes were chosen phenomenologically to produce qualitatively distinct dynamics; empirical quantification of stress-induced switching rates in tumor cells under acidic conditions remains an important area for future experimental work.
@@
Global Sensitivity Analysis: To quantify parameter influence on equilibrium resistance while accounting for parameter interactions, we performed global sensitivity analysis using Latin Hypercube Sampling (LHS) with Partial Rank Correlation Coefficients (PRCC) [27]. We generated $N=1000$ samples across the 17-dimensional parameter space, with each parameter sampled uniformly from its biologically plausible range (Table 1). For each sample, the system was integrated to equilibrium and the resistant fraction $x^{*}$ recorded. PRCC values quantify the monotonic relationship between each parameter and $x^{*}$ while controlling for all other parameters.
@@ 3.2.3 Bifurcation with Respect to Protection Factor
Critical clearance ($\gamma_{0}=0.039$ day-1): At the critical clearance rate, protection factor produces a sharp bifurcation. Above $\varphi\approx 0.58$, the system equilibrates to high resistance ($x^{*}>0.95$), while below this threshold, resistance is almost suppressed ($x^{*}<0.1$). This indicates that at marginal perfusion levels, acid-resistance machinery effectiveness becomes a critical determinant of evolutionary outcomes.
@@
Figure 3: Regime A: Context-dependent bifurcation with respect to protection factor. Equilibrium resistant fraction $x^{*}$ as a function of protection factor, $\varphi$, at three clearance rates: $\gamma_{0}=0.025$ day-1 (below bifurcation threshold), $\gamma_{0}=0.039$ day-1 (at bifurcation threshold), and $\gamma_{0}=0.052$ day-1 (above bifurcation threshold). Protection factor produces a sharp bifurcation only at intermediate clearance rates, demonstrating parameter interaction. At low clearance, resistance is inevitable regardless of $\varphi$; at high clearance, sensitivity is maintained regardless of $\varphi$. This indicates that targeting acid-resistance machinery is only effective within a specific range of vascular perfusion levels.
@@ 3.3.3 Combined Effects of Clearance and Protection
3.4 Sensitivity Analysis
[…]
Figure 7: Global sensitivity analysis of equilibrium resistance. Partial Rank Correlation Coefficients (PRCC) quantifying parameter influence on equilibrium resistant fraction $x^{*}$, computed from $N=1000$ Latin Hypercube samples. Positive values (red) indicate parameters whose increase promotes resistance; negative values (blue) indicate suppression of resistance; gray indicates no significant correlation ($p>0.05$). The most influential parameters are $h_{c}$, $\mu_{0}$, $\gamma_{0}$, $c$, $\alpha$, and $r$. Model predictions are robust to $\beta$ and $\eta$.
@@ 4 Discussion and Conclusions
Several limitations of the present framework warrant discussion. First, the spatially homogeneous assumption neglects pH gradients and localized selection pressures that arise in heterogeneous tumor microenvironments; spatial extensions using reaction-diffusion formulations could capture these dynamics [34, 35]. Second, the binary phenotype classification simplifies a likely continuous spectrum of acid resistance; partial differential equation models structured by resistance level represent a natural generalization [36]. Third, the model omits immune cell populations, whose function is known to be impaired under acidic conditions [7, 10]; incorporating tumor-immune interactions could reveal additional therapeutic opportunities. Fourth, vasculature is treated as static, whereas dynamic angiogenesis and vascular remodeling influence both acid clearance and tumor growth [3, 28]; coupling to angiogenesis models would enhance physiological realism. Finally, the absence of clinical or experimental validation limits predictive confidence; fitting to longitudinal measurements of tumor pH and resistance markers represents an essential next step [1].
Our model’s identification of plasticity as a critical determinant of treatment response emphasizes the need for biomarkers distinguishing these tumor phenotypes. Potential markers of phenotypic plasticity include expression levels of epigenetic modifiers (EZH2, DNMT1), epithelial-mesenchymal transition markers, or functional assays measuring adaptation rates to environmental stress [14, 15, salgia2018plasticity]. methods to assess tumor phenotypic plasticity. Unlike conventional biomarkers that capture static tumor state at a single timepoint, plasticity is a dynamic property representing the rate at which cells switch phenotypes under stress. Estimating the switching rate $\mu_{0}$ would require functional assays that expose tumor cells to controlled acidic conditions and measure the rate of phenotypic adaptation over time, for instance by tracking the emergence of acid-resistant subpopulations via flow cytometry or live-cell imaging [14, 15]. Developing such assays for clinical use represents a necessary step toward implementing regime-dependent therapeutic strategies. This framework provides a foundation for regime-dependent interventions where treatment strategies are tailored to the tumor’s evolutionary capacity: perfusion-focused therapy for low-plasticity tumors versus combination approaches for high-plasticity tumors.